# Amandine Brunetto  Sascha Hornauer  Fabien Moutarde Centre for Robotics

### v1

@@ 4.2 Neural Radiance Field
NeRAF relies on Nerfacto [53] as its NeRF model. Nerfacto is a combination of many works following [34] including camera pose refinement [57, 29], per image appearance conditioning [33], proposal sampling [35], scene contraction and hash encoding. We chose this approach because of its strong performance on real-world data captures of static scenes. However, our method is architecture-agnostic, as it only relies on the radiance field and does not modify the NeRF.
4.3 Grid Sampler
Figure 3: Grid sampler. We represent the scene as a grid of voxels. The grid is filled by querying the radiance field with voxel center coordinates $\textbf{X}_{vi}$ and multiple viewing directions $\textbf{d}_{1\rightarrow 18}$. We average the color values obtained for each viewing direction. It results in a 7-channels 3D grid containing color $\mathbf{\widehat{C}}_{vi}$, density $\sigma_{vi}$ and voxel-centers 3D coordinates.
Sound propagation is omni-directional and determined by 3D geometry and material properties of an environment, thus we provide 3D information about the scene to our neural acoustic field.
To this end, we use our grid sampler, shown in Figure 3, to construct a 3D volume from NeRF. Specifically, we represent the scene by a voxel grid of size $[0,1]^{3}$ with a $1/gap$ resolution. We apply the same scene contraction as NeRF.
We obtain the positions of voxel centers $\textbf{X}_{vi}$ and update the grid at each training iteration using a batch of $4,096$ $\textbf{X}_{vi}$. This sampling approach ensures that the grid is fully populated, unlike directly using the $\textbf{X}_{n}$ sampled by NeRF proposal sampling. Indeed, NeRF learns to sample $\textbf{X}_{n}$ likely to be of interest, which could potentially leave some regions of the grid unfilled or without further updates as the model improves.
As NeRF color is view-dependent, we follow [12] and form $N=18$ viewing rays per $\textbf{X}_{vi}$. The radiance fields return a density value $\sigma_{vi}$ and a color value $\textbf{c}_{vi_{j}}$ per $j\in[0,N]$ viewing directions:
$\text{NeRF}:(\textbf{X}_{vi},\textbf{d}_{1\rightarrow N})\rightarrow(\textbf{c}_{vi_{1\rightarrow N},\sigma_{vi}})$ (4)
For color, we use the volume rendering equation and average values over all viewing directions ĉ$_{vi}$. For density, we compute the alpha compositing value $\alpha=1-\exp(-\sigma_{vi}\delta)$ where $\delta$ is a chosen small value. We fill the grid with these values, resulting in 4-channels, and concatenate the $\textbf{X}_{vi}$ 3D coordinates to add spatial information which is important for sound propagation.
4.4 Neural Acoustic Field
The goal of our neural acoustic field (NAcF), presented in Figure 4, is to learn a continuous neural representation of the scene’s acoustics. To this end, NAcF maps microphone coordinates $\textbf{X}_{m}=(x_{m},y_{m},z_{m})$ and orientations $\textbf{d}_{m}=(\theta_{m},\phi_{m})$ along with source position $\textbf{X}_{s}=(x_{s},y_{s},z_{s})$ and orientation $\textbf{d}_{s}=(\theta_{s},\phi_{s})$ to a binaural room impulse response:
$\text{NAcF}:(\textbf{X}_{m},\textbf{d}_{m},\textbf{X}_{s},\textbf{d}_{s},t)\rightarrow\text{RIR}(f,t)$ (5)
Figure 4: Neural Acoustic Field. NAcF maps microphone-source poses and directions to binaural RIRs. It is conditioned by scene features obtained via a 3D ResNet. Poses, directions and time queries are projected to a higher dimensional space through positional encoding. NAcF outputs a vector containing $F$ frequency values for each time query.
Soundspaces dataset lacks the complete acoustic field parameterization described in Equation 5. Microphone rotates only around the up-axis, $\theta$, and the source is omni-directional. Thus, NAcF is parametrized only by $(\textbf{X}_{m},\textbf{d}_{m}=\theta,\textbf{X}_{s},t)$.
We encode audio using short-time Fourier transform (STFT) resulting in $\text{STFT}\in\mathbb{R}^{C\times F\times T}$ with $C$, $F$ and $T$ respectively the channels, frequency and time bins. Compared to time domain, the time-frequency representation is smoother and more suitable to neural network prediction [30]. NAcF learns the log-magnitude of the STFT. We query time $t$ because it is particularly important for RIR: it’s crucial to perceptual effects such as reverberation. NAcF returns the frequency values corresponding to the time query. We encode $t$ between $[-1,1]$ and use positional encoding. Similar to NeRF [34] pose processing, microphone and source coordinates as well as direction are also embedded to a higher dimensional space using positional encoding.
Given the voxel grid, we concurrently train a ResNet3D [25] to extract pertinent features for sound propagation. They serve as implicit indicators of scene geometry and material properties which are central for acoustic modeling. This choice follows works [11, 47, 28] in audio-visual research showing that extracted geometry and material information using ResNet networks contribute to audio synthesis. We perform multimodal fusion by concatenating grid features vector with the encoded microphone and source poses, microphone direction and time query. Similar to NeRF, NAcF consists of two MLP blocks. The first block maps the input to a feature vector embedding the input information and the acoustic scene. The second learns to spatialize sound, i.e. the head related transfer function (HRTF) in the case of binaural microphones. It contains one MLP per microphone channel and output the $F=(N_{fft}/2)+1$ frequencies for each time query.
Contrary to previous works, NAcF does not directly predict the complete STFT during training. Instead, we adopt a process similar to NeRF’s. NeRF shuffles all image pixels present in the train set and creates batches independently of their images of origin. In NeRAF, we shuffle all STFT time bins in the train set and form batches, independently of their STFT of origin. This method efficiently trains the network to learn how sound propagates over time. At inference, we query all time bins $t\in[0,T]$ to render the complete STFT.
Finally, we use Griffin-Lim [23, 38] algorithm to obtain RIR waveforms from magnitude STFTs.
4.5 Learning Objective
Cross-modal learning has proven beneficial in multiple works. Thus, NeRAF trains jointly NeRF and NaCF. NAcF training is slightly delayed as it begins when the grid has been updated a few times.
We use NeRF MSE loss function $\mathcal{L}_{V}$ between ground-truth pixel color $\hat{C}(\textbf{r})$ and the rendered one $C(\textbf{r)}$) :
[…]
Note that each NeRF method also has complementary losses.
@@ 4.5 Learning Objective
Inspired by [59, 60], we use a combination of spectral loss $\mathcal{L}_{SL}$ [16] and spectral convergence loss $\mathcal{L}_{SC}$ [1] as our audio loss $\mathcal{L}_{A}$:
$\mathcal{L}_{A}=\lambda_{SC}\mathcal{L}_{SC}+\lambda_{SL}\mathcal{L}_{SL}$ (7)
$\mathcal{L}_{SL}=\lVert\log(\left|\hat{s}\right|+\epsilon)-\log(\left|s\right|+\epsilon)\rVert_{2}^{2}$ (8)
$\mathcal{L}_{SC}=\frac{\lVert\left|\hat{s}\right|-\left|s\right|\rVert_{F}}{\lVert\left|\hat{s}\right|\rVert_{F}}$ (9)
@@
where $\hat{s}$ is the ground-truth STFT, $s$ the predicted STFT, .$\lVert.\rVert_{F}$ denotes the Frobenius norm, .$\lVert.\rVert_{2}$ the L2 norm and $\epsilon=10^{-3}$. The spectral convergence loss emphasizes on spectral peaks helping especially in early phases of training while the spectral loss focus on small amplitude components which tends to be more important towards the later phases of training. Our loss differs from [59, 60] as we found out that MSE spectral loss leads to better performance than L1 spectral loss. Additionally, our loss combination outperforms the MSE loss on magnitude STFT used in [30, 28]. The loss comparison is presented in Section 5.3.
NeRAF complete learning objective can be resumed as:
$\mathcal{L}=\mathcal{L}_{V}+\lambda_{A}\mathcal{L}_{A}$ (10)
@@ 5 Experiments
First, we evaluate NeRAF’s performance against NAF, INRAS and AV-NeRF on the SoundSpaces dataset. Then, we demonstrate that cross-modal learning of the acoustic and radiance fields is beneficial to vision, resulting in enhanced results for large complex scenes with very sparse training observations. Next, we conduct a few-shot experiment to assess our method with a reduced number of training audio recordings. Finally, we investigate the impact of the 3D voxel-grid and of our loss combination on NeRAF.
@@ 5.1 Datasets
To the best of our knowledge, there is no existing real-world dataset that includes RGB images with camera poses and orientations, along with RIRs and microphone-source poses. Especially, for NeRF to perform well, training views must sufficiently cover the scene. Therefore, we follow previous works and evaluate NeRAF on the SoundSpaces simulated dataset.
@@
SoundSpaces [6] is an audio simulator built upon Habitat Sim [45, 52, 39], a 3D simulator for embodied AI research. It provides binaural RIRs at discrete positions of a 2D grid with a spatial resolution of $0.5$ m with four head orientations accompanied by sound source position. This popular framework [21, 7, 30, 28, 51, 40, 36, 8] enables us to evaluate our method on diverse scene types while being under same settings as previous works. Following [30, 28, 51] we select six indoor scenes from Replica [50]. Two are small rooms (office 4 and room 2) with rectangular walls, two are medium room (frl apartment 2 and 4) with more complex layout and objects and two are complete apartments containing multiple rooms (apartment 1 and 2). For each scene, we use SoundSpaces pre-generated binaural RIRs with corresponding source and microphone positions. To train NeRF, we generate RGB images using Habitat Sim. Camera poses and parameters were selected to limit the amount of visual data. More details on the motivations and on the generation process are available in the supplementary material.
5.2 Evaluation
Table 1: Comparison with State-of-the-art. Performance on the SoundSpaces dataset using T60, C50, and EDT metrics. Lower score indicates higher RIR quality.
Methods T60 Error (%) $\downarrow$ C50 Error (dB) $\downarrow$ EDT Error (sec) $\downarrow$
@@ 5.2 Evaluation
NAF [30] 3.18 1.06 0.031
INRAS [51] 3.14 0.60 0.019
AV-NeRF [28] 2.47 0.57 0.016
NeRAF 2.04 0.39 0.011
@@
To assess the effect of cross-modal learning on novel view synthesis, we employ commonly used metrics: PSNR, SSIM, and LPIPS.
@@
Similar to [51, 30, 28], we use $90\%$ of SoundSpaces audio data for training and $10\%$ for testing. We resample the RIR to $22,050$ Hz and compute STFT using $N_{fft}=512$ and $\text{{hop}}=128$. To train the vision part, we generate images using SoundSpaces 2.0 [9]. For small, medium and large rooms, we use respectively $45$, $75$ and $150$ observations. Evaluation set comprises 50 images.
@@ Results.
We compare NeRAF with state-of-the-arts methods on SoundSpaces in Table 1. Similar to previous works on neural acoustic field, we also compare our method to traditional audio encoding methods AAC and Opus [17, 18]. AAC is a multichannel audio coding standard and Opus is an open audio codec. With a grid of resolution $128$, NeRAF significantly outruns existing methods. It achieves $17.4\%$ improvement on T60, $31.3\%$ on EDT and $31.6\%$ on C50 compared to AV-NeRF. Since they also employ Nerfacto, our performance improvement is not due to a superior NeRF architecture, but rather to our method’s design. We showcase qualitative results of audio synthesis in supplementary material.
@@
We next explore the effect of jointly learning acoustics and radiance fields on complex scenes with a low amount of training views. We train NeRAF and Nerfacto on the two large rooms. 75, 100, 150 images are used for training and the same 50 novel views are used for evaluation. In Table 2 we show that joint training of acoustic and radiance field improves novel view synthesis. It has a stronger impact on LPIPS which is related to the human-perceived similarity between two images. Overall, the performance improvement is more significant when the training images are very sparse. We present qualitative results in Figure 5.
Table 2: Quantitative Results on Cross-Modal Learning. For highly complex scenes with very sparse training images, NeRAF audio-visual joint training improves vision performances, especially LPIPS. Evaluation on 50 test novel views. Delta are computed against Nerfacto (no sound). Metrics are averaged over the two large rooms.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Training Images 75 100 150 75 100 150 75 100 150
$\Delta\text{NeRAF}$ +3.62% +2.24% +1.23% +1.44% +1.53% +1.03% -17.11% -9.86% -6.25%
@@
Previous works rely on a significant amount of RIR data, which is often unavailable in real-life scenarios. Therefore, we assess the performance of our model when trained with fewer data. We conduct a few-shot experiment on one small, one medium, and one large room and average the results. As shown in Figure 6, with only half the data, NeRAF outperforms AV-NeRF trained on the complete training set.
@@ 5.3 Ablation Study
We compare NeRAF performances without grid and with grids of resolution $128^{3}$ and $256^{3}$. Results are presented in Table 3 (Left). Without the voxel grid, NAcF is no longer conditioned by the 3D information from the scene. It still outperforms AV-NeRF with $10.5\%$, $28.1\%$ and $31.3\%$ improvement on T60 error, C50 and EDT. Compared to no grid, the 128-resolution 3D grid improves T60 error by $7.5\%$ and C50 by $5.3\%$. We do not observe an enhancement in EDT. However, increasing the resolution to 256 improves EDT by $5.9\%$. EDT is related to RIR early reflections which usually corresponds to obstacles very close to the listener. Thus, it makes sense that the method needs greater resolution i.e. smaller voxels to perceive those obstacles. Increasing the grid resolution also improves C50, but we observe a decrease in T60 performance. T60 is related to the late reverberation of the RIR, which depends mainly on the size of the room. Consequently, increasing grid resolution should not significantly improve the quality of this information.
Figure 6: Few-shot Learning. Average on office 4, frl apartment 2 and apartment 2. With just half the data, NeRAF outperforms AV-NeRF trained on all the training data.
Table 3: Ablation study. Bold and underlined respectively indicates best and second best. Left: Influence of the grid. Evaluation on all scenes in SoundSpaces dataset. NG stands for no grid, while 128 and 256 respectively correspond to grid resolutions $128^{3}$ and $256^{3}$. Right: Loss study. Performances averaged across office 4, frl apartment 2 and apartment 2.
Methods T60 (%) $\downarrow$ C50 (dB) $\downarrow$ EDT (sec) $\downarrow$
NG 2.21 0.41 $0.011_{4}$
128 2.04 0.39 0.0114
256 2.38 0.388 0.0107
Methods T60 (%) $\downarrow$ C50 (dB) $\downarrow$ EDT (sec) $\downarrow$
MSE 2.20 0.52 0.014
SC+SLL1 1.98 0.57 0.016
Ours 2.15 0.37 0.010
@@
We examine the impact of our custom loss described in Section 4.5 compared to the MSE loss used in previous works [28, 30] and the combination with L1 spectral loss proposed by [60]. To this end, we evaluate NeRAF’s performance using these losses. As shown in Table 3 (Right), compared to the MSE loss, our combination significantly improves EDT and C50 by 29.1% and 27.9%, respectively, and slightly enhances T60 by 2.21%. Additionally, using MSE for $\mathcal{L}_{SL}$ instead of L1 substantially improves C50 and EDT by 34.1% and 36.4%, although it results in 9% decline in T60.

### v2

@@ 3.5 Learning Objective
Inspired by (Yamamoto et al., 2019; 2020), we use a combination of spectral loss $\mathcal{L}_{\text{SL}}$ (Défossez et al., 2018) and spectral convergence loss $\mathcal{L}_{\text{SC}}$ (Arık et al., 2018) as our audio loss $\mathcal{L}_{A}$:
$\mathcal{L}_{A}=\lambda_{\text{SC}}\mathcal{L}_{\text{SC}}+\lambda_{\text{SL}}\mathcal{L}_{\text{SL}},$ (7)
$\mathcal{L}_{\text{SL}}=\lVert\log(\left|s\right|+\epsilon)-\log(\left|\hat{s}\right|+\epsilon)\rVert_{2}^{2},$ (8)
$\mathcal{L}_{\text{SC}}=\frac{\lVert\left|s\right|-\left|\hat{s}\right|\rVert_{F}}{\lVert\left|s\right|\rVert_{F}},$ (9)
@@
where $\hat{s}$ is the predicted STFT, $s$ the ground-truth STFT, .$\lVert.\rVert_{F}$ denotes the Frobenius norm, .$\lVert.\rVert_{2}$ the L2 norm and $\epsilon=10^{-3}$. The values of our loss hyperparameters are found empirically and given in Appendix C. The spectral convergence loss emphasizes on spectral peaks helping especially in early phases of training while the spectral loss focus on small amplitude components which tends to be more important towards the later phases of training. Our loss differs from (Yamamoto et al., 2019; 2020) as we found out that the MSE spectral loss leads to a better trade-off between metrics than the L1 spectral loss (Su et al., 2022) and the MSE loss on magnitude STFT (Luo et al., 2022; Liang et al., 2023a). NeRAF results with those three losses are presented in Section 4.3. The complete learning objective can be resumed as:
$\mathcal{L}=\mathcal{L}_{V}+\lambda_{A}\mathcal{L}_{A}.$ (10)
@@ 4 Experiments
Figure 5: Loudness maps visualization. We visualize the intensity of predicted RIRs at different microphone positions for a given loudspeaker position and orientation. Intensities are averaged over multiple heights.
First, we evaluate NeRAF’s performance against related works (Luo et al., 2022; Su et al., 2022; Liang et al., 2023a; b) on SoundSpaces and RAF datasets. Then, we demonstrate that cross-modal learning of the acoustic and radiance fields is also beneficial to vision, resulting in enhanced results for large complex scenes trained with very sparse observations. Next, we conduct a few-shot experiment to assess the effectiveness of our method when trained with a limited number of audio recordings. Finally, we investigate the impact of the 3D grid and our loss combination on NeRAF’s performance. Video examples showing NeRAF audio-visual generation on real and synthetic scenes are presented in the supplementary material.
@@ 4.1 Datasets
SoundSpaces (Chen et al., 2020a) is an audio simulator built upon Habitat Sim (Savva et al., 2019; Szot et al., 2021; Puig et al., 2024), a 3D simulator for embodied AI research. It provides binaural RIRs at discrete positions of a 2D grid with a spatial resolution of $0.5$ m with four head orientations accompanied by sound source position. This popular benchmark (Gao et al., 2020; Chen et al., 2020b; Luo et al., 2022; Liang et al., 2023a; Su et al., 2022; Purushwalkam et al., 2021; Parida et al., 2021; Chen et al., 2022a) enables us to evaluate our method on diverse scene types while being under same settings as previous works. Following (Luo et al., 2022; Liang et al., 2023a; Su et al., 2022; Liang et al., 2023b), we select six indoor scenes from Replica (Straub et al., 2019). Two are small rooms (office 4 and room 2) with rectangular walls, two are medium room (frl apartment 2 and 4) with more complex layout and objects and two are complete apartments containing multiple rooms (apartment 1 and 2).
RAF.
Real Acoustic Fields (RAF) (Chen et al., 2024) is a real-world dataset that includes high-quality, densely captured room impulse response data paired with multi-view images and precise 6DoF pose tracking data for sound emitters and listeners in two rooms. The room impulse response are recorded using the ”earful tower”, which is equipped with omnidirectional microphones. This dataset was recently used to benchmark previous works on real-world scenes.
RWAVS.
RWAVS (Liang et al., 2023a) is a real-world dataset used to benchmark AV-NeRF. It does not contain RIRs but audio recordings. Our results on the RWAVS dataset are presented in Appendix J.
Table 1: Comparison with state-of-the-art. Performance on the SoundSpaces dataset using T60, C50, and EDT metrics (Left) and on RAF (Right) with the same metrics and STFT error. Lower score indicates higher RIR quality.
Methods T60 (%) $\downarrow$ C50 (dB) $\downarrow$ EDT (sec) $\downarrow$
@@
NAF 3.18 1.06 0.031
INRAS 3.14 0.60 0.019
AV-NeRF 2.47 0.57 0.016
NACF 2.36 0.50 0.014
NACF w/ T 2.17 0.49 0.014
NeRAF 2.14 0.38 0.010
Methods T60 (%) $\downarrow$ C50 (dB) $\downarrow$ EDT (sec) $\downarrow$ STFT error (dB) $\downarrow$
NAF 10.08 0.71 0.021 0.64
INRAS 8.01 0.79 0.025 0.36
AV-NeRF 8.11 0.73 0.021 0.39
NeRAF 7.47 0.61 0.020 0.17
NAF++ 8.19 0.53 0.017 0.64
INRAS++ 6.17 0.57 0.017 0.39
NACF 6.62 0.59 0.017 0.39
NACF w/ T 7.31 0.59 0.018 0.39
NeRAF++ 6.87 0.51 0.015 0.17
4.2 Evaluation
@@
Similar to (Su et al., 2022; Luo et al., 2022; Liang et al., 2023a), we use $90\%$ of SoundSpaces audio data for training and $10\%$ for testing. They are pre-generated binaural RIRs with corresponding source and microphone positions. We resample the RIR to $22.05$ kHz and compute STFT using $N_{\text{FFT}}=W=512$ and $H=128$ where $N_{\text{FFT}}$ is the number of FFT bins, W the Hann window size and H the hop length. To train the vision part, we generate RGB images using Habitat Sim. Camera poses and parameters were selected to limit the amount of visual data. For small, medium and large rooms, we use respectively $45$, $75$ and $150$ observations. Evaluation set comprises 50 images. More details on the motivations and the generation process are available in Appendix D.
For RAF, we follow previous works experimental setup: we keep $80\%$ of data for training and $20\%$ for evaluation. Audios are cut to $0.32$ s, resampled to $48$ kHz and processed into STFT using $N_{\text{FFT}}=1024$, $W=512$ and $H=256$. For the vision part, we randomly select one third of the data from cameras 20 to 25. Nerfstudio automatically keep 90% of them for training and 10% for evaluation.
To ensure statistically significant results, we conduct each experiment 6 times and average results. More details on the implementation and architecture are presented in Appendix C, Appendix B.
Figure 7: Few-shot Learning. With 25% less audio recordings, NeRAF outperforms AV-NeRF trained on all the training data. Results are averaged on office 4, frl apartment 2 and apartment 2.
@@ Results.
We compare NeRAF with state-of-the-arts methods on SoundSpaces in Table 1 (Left). Similar to previous works on neural acoustic field, we also compare our method to traditional audio encoding methods AAC and Opus (for Standardization, 2006; Foundation, 2012). AAC is a multichannel audio coding standard and Opus is an open audio codec. NeRAF significantly outruns existing methods for C50 and EDT. It achieves $22.4\%$ improvement on C50 and $28.6\%$ on EDT compared to NACF with temporal alignment. Compared to AV-NeRF – which is the method most similar to ours as it also employ Nerfacto and does not rely on a large amount of information known a priori – we achieve respectively $13.4\%$, $33.3\%$ and $37.5\%$ improvement for T60, C50 and EDT.
As SoundSpaces is a simulated dataset that presents limitations in term of visual and acoustics realism, we also present results on RAF in Table 1 (Right). NeRAF outperforms previous works best results, achieving $-6.7$% T60, $-14.2$% C50, $-4.5$% EDT and $-52.5$% STFT error. The second part of the table contains a comparison of methods using an additional energy decay loss term (Majumder et al., 2022; Liang et al., 2023b). While NACF initially relies on this term, other methods, with this new term, are denoted by ”++”. NeRAF++ still achieves state-of-the-art results, in particular for C50, EDT and STFT error. We present details about this loss in Appendix C. We showcase qualitative results of audio synthesis in Figure 5 and in Appendix I.
Table 2: Quantitative results on cross-modal learning. For highly complex scenes with very sparse training images, NeRAF audio-visual joint training improves vision performances, especially LPIPS. Evaluation on 50 test novel views. Delta are computed against Nerfacto (no audio).
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
# Images 75 100 150 75 100 150 75 100 150
Apartment 1
Nerfacto 22.48 24.77 27.00 0.838 0.882 0.902 0.246 0.161 0.118
NeRAF 23.33 25.07 27.70 0.858 0.884 0.915 0.216 0.140 0.105
$\Delta\text{NeRAF}$ +3.78% +1.23% +2.60% +2.39% +0.28% +1.37% -12.20% -13.22% -11.25%
Apartment 2
Nerfacto 19.10 22.67 25.62 0.777 0.836 0.884 0.381 0.232 0.160
NeRAF 20.53 23.46 26.06 0.802 0.856 0.895 0.322 0.206 0.150
$\Delta\text{NeRAF}$ +7.48% +3.49% +1.69% +3.26% +2.31% +1.20% -15.27% -11.14% -6.03%
@@
We next investigate the impact of jointly learning acoustic and radiance fields on vision in complex scenes with very sparse training views, a challenging scenario where NeRF typically struggles. We train NeRAF and Nerfacto on the two large rooms. 75, 100, 150 images are used for training and the same 50 novel views are used for evaluation. In Table 2 we show that joint training of cross-modal learning improves novel view synthesis. It has a stronger impact on LPIPS which is related to the human-perceived similarity between two images. Overall, the performance improvement is more significant when the training images are very sparse. We present qualitative results in Figure 6: NeRAF reduces floating artifacts and enables more detailed images. Results on real scenes and additional visualizations are presented respectively in Appendix J and Appendix I.
@@
Previous works evaluated on SoundSpaces rely on a significant amount of RIR data, which is often unavailable in real-life scenarios. Therefore, we assess the performance of our model when trained with fewer data. We conduct a few-shot experiment on one small, one medium, and one large room and average the results. As shown in Figure 7, NeRAF is better than AV-NeRF on all metrics when trained with 75% of the data. It still outperforms AV-NeRF on EDT when trained with only 25% of the data. On RAF dataset – which is a real-world dataset that already contains a more realistic amount of training data – NeRAF still outperforms the previous methods when trained with 25% less data. Results on RAF are presented in Appendix E.
@@ 4.3 Ablation Study
We assess the grid impact on performances in Table 3 (Left) by first removing the grid (NG) and then filling the 128-grid with zeros (128 w/ 0). In both cases, NAcF is no longer conditioned by the 3D priors from the scene. Removing the grid also means we reduce the number of trainable parameters as we no longer keep the features extractor. By setting the grid to zero we remove scene-related information and still keep the exact same architecture. This way, we show that the performance gap cannot be explained only by an increased number of parameters but rather to the information extracted from the 3D-grid. We also study the impact of the resolution of the grid by comparing $S=128$ and $S=256$. Increasing the resolution to 256 improves T60 and slightly C50 but is accompanied by a minor EDT decline and increased training computation time.
Table 3: Ablation studies. Left: Influence of the grid evaluated on RAF dataset. NG stands for no grid, 128 w/ 0 for a 128-grid filled with zeros, while 128 and 256 respectively correspond to grid of resolutions $S=128$ and $S=256$. Bold and underlined respectively indicates best and second best. Right: NeRAF’s loss study. Performances averaged across SoundSpaces dataset.
Grid T60 (%) $\downarrow$ C50 (dB) $\downarrow$ EDT (sec) $\downarrow$
NG 8.15 0.674 0.0212
128 w/ 0 7.75 0.626 0.0205
128 7.47 0.609 0.0201
256 7.42 0.608 0.0204
Loss T60 (%) $\downarrow$ C50 (dB) $\downarrow$ EDT (sec) $\downarrow$
MSE 1.89 0.53 0.015
[…]
@@
[…]